# A Reduction-Driven Local Search for the Generalized Independent Set Problem

### v1

@@ 6.3 Performance Comparison with Existing Algorithms
Instance $|V|$ $|V_{ker}|$ LA-B&B RLS
without reduction with reduction without reduction with reduction
BKV time(s) BKV time(s) BKV htime(s) BKV htime(s)
bio-yeast 1,458 0 66,545 300 68,574 0 68,574 1 68,574 $<$1
soc-wiki-Vote 889 0 35,973 300 37,358 0 37,358 1 37,358 $<$1
vc-exact_131 2,980 0 107,812 300 109,641 0 109,567 9 109,641 $<$1
web-edu 3,031 98 109,856 300 113,781 15 114,080 12 114,140 $<$1
MANN_a81 3,321 0 114,616 300 117,327 0 117,275 14 117,327 $<$1
tech-routers-rf 2,113 0 93,690 300 97,799 0 97,799 1 97,799 $<$1
vc-exact_039 6,795 0 254,941 300 259,542 0 259,475 13 259,542 $<$1
ca-GrQc 4,158 282 167,846 300 174,654 15 174,768 7 174,782 11
vc-exact_038 786 697 10,371 300 12,014 12 12,015 13 12,015 4
vc-exact_078 11,349 0 102,312 300 438,671 1 438,306 16 438,671 $<$1
vc-exact_087 13,590 0 71,804 300 518,500 0 518,069 15 518,500 $<$1
vc-exact_107 6,402 0 88,602 300 407,878 0 406,379 14 407,878 $<$1
vc-exact_151 15,783 0 54,114 300 530,138 0 529,003 18 530,138 $<$1
vc-exact_167 15,783 0 55,190 300 465,178 0 463,446 12 465,178 $<$1
bio-dmela 7,393 44 280,897 300 302,449 10 302,793 18 302,992 $<$1
vc-exact_011 9,877 172 190,729 300 301789 7 303,352 9 303,793 $<$1
web-spam 4,767 487 180,446 300 190,344 14 190,460 18 190,459 13
vc-exact_026 6,140 300 177,898 300 201,316 17 201,458 13 201,503 $<$1
vc-exact_001 6,160 134 190,729 300 215,668 13 215,934 13 215,941 $<$1
vc-exact_024 7,620 1,255 226,540 300 235,440 14 235,911 14 236,018 12
C1000.9 1,000 1,000 6,102 300 6,102 300 7,235 16 7,237 10
tech-WHOIS 7,476 338 309,903 300 325,327 15 325,480 16 325,581 13
p-hat700-3 700 700 2,937 300 2,937 300 4,473 1 4,473 $<$1
vc-exact_008 7,537 574 71,804 300 246,771 14 247,036 16 247,091 $<$1
keller5 776 776 2,788 300 2,788 300 3,877 2 3,877 2
vc-exact_194 1,150 1,150 4,273 300 4,273 300 5419 2 5,419 $<$1
hamming10-4 1,024 1,024 2,871 300 2,871 300 3660 1 3660 $<$1
brock800_2 800 800 1,996 300 1,996 300 2473 1 2473 $<$1
brock800_4 800 800 1,968 300 1,968 300 2501 1 2501 $<$1
ca-HepPh 11,204 2,165 153,907 300 372,079 6 372,079 14 372,079 17
p-hat700-2 700 700 2,127 300 2,127 300 2,971 1 2,971 $<$1
vc-exact_196 1,534 1,534 4,273 300 4,273 300 5,321 2 5,321 $<$1
p-hat700-1 700 700 844 124 844 124 844 1 844 $<$1
ca-AstroPh 17,903 4,985 44,296 300 602,320 2 603,732 15 603,945 14
C2000.9 2,000 2,000 5,477 300 5,477 300 6,396 3 6,396 3
DSJC1000.5 1,000 1,000 1,492 300 1,492 300 1,861 1 1,861 $<$1
socfb-MIT 6,402 5,494 110,677 300 129,807 17 129,831 16 129,872 16
p-hat1500-3 1,500 1,500 3,856 300 3,856 300 7,500 15 7,501 16
socfb-UCSB37 14,917 13,655 219,965 300 254,965 16 255,887 21 255,898 22
socfb-Duke14 9,885 9,003 125,940 300 151,169 15 151,321 14 151,326 18
p-hat1500-2 1,500 1,500 2444 300 2,444 300 4,253 1 4,253 $<$1
socfb-Stanford3 11,586 9,380 215,381 300 251,350 12 251,788 19 251,833 14
socfb-UConn 17,206 16,025 170,548 300 274,637 12 275,006 16 275,034 20
p-hat1500-1 1500 1,500 1,121 300 1,121 300 1,287 2 1,287 1
C2000.5 2,000 2,000 1,570 300 1,570 300 1,849 5 1,849 6
keller6 3,361 3,361 3,545 300 3,545 300 5,229 19 5,236 16
C4000.5 4,000 4,000 1,601 300 1,601 300 1,951 15 1,951 22
foodweb-wet 128 125 2,464 55 2,464 27 2,464 $<$1 2,464 $<$1
foodweb-dry 128 125 2,226 28 2,226 25 2,226 $<$1 2,226 $<$1
USAir97 332 0 9,507 2000 10,308 0 10,308 $<$1 10,308 $<$1
powergrid 4,941 0 190,537 2000 193,339 0 193,194 52 193,339 $<$1
CondMat 23,133 1,648 95,130 2000 708,907 2000 709,927 54 711,334 45
Email 265,009 0 OOM 13,054,427 1 13,054,390 69 13,054,427 1
Epinion 75,879 253 7,400 2000 3,051,256 2000 3,050,415 67 3,052,059 1
Dblp 317,080 11,851 OOM 9,715,860 2000 9,577,609 46 9,734,653 64
cnr-2000 325,557 59,052 OOM 11,347,241 2000 13,361,167 84 13,368,670 482
WikiTalk 2,394,385 8 OOM 118,708,326 $<$1 118,672,055 56 118,708,326 333
BerkStan 685,230 114,191 OOM 22,591,394 2000 25,813,846 2000 26,031,704 1112
As-Skitter 1,696,415 31,173 OOM 65,907,243 2000 59,886,304 2000 66,582,254 723
soc-pokec 1,632,803 881,319 OOM OOM 36,540,838 350 43,999,402 934
LiveJ 4,846,609 401,307 OOM OOM 148,758,359 318 168,431,056 904
uk2002 18,483,186 4,806,660 OOM OOM 633,365,020 1806 687,089,901 7678
# of best / # of instances 3/62 24/62 24/62 62/62
average running time(s) 415.2 296.7 119.4 202.6
Table 4 compares algorithms on SET-3 instances. OOM indicates that the algorithm runs out of memory. Both ALS and LA-B&B quickly become infeasible as instance size increases. In contrast, RLS computes feasible solutions even for large instances such as uk2002 . For instances where the benchmarks are feasible, RLS consistently achieves the best net benefit. For the infeasible instance Email, RLS even finds the optimal solution.
6.4 Evaluation of Reduction Rules
To evaluate the effectiveness of the reduction rules, we conduct the following two experiments.
Comparative results with vs. without reduction. This experiment evaluates algorithms with and without data reduction. For LA-B&B with reduction, we use kernel graphs as input. For RLS without reduction, all reductions are skipped, and random initialization is applied. ALS cannot handle kernel graphs due to unsupported negative vertex profits and edge penalties.
The comparative results are shown in Table 5. In general, the reduction rules can help both algorithms to find better solutions. Specifically, with the reduction technique, LA-B&B is able to solve some instances which it cannot solve without reduction, e.g. WikiTalk (1.6 million vertices, 11.1 million edges). And for some large instances that LA-B&B fails due to out-of-memory, it is able to report a feasible solution or even solve them. As for RLS, with these reduction rules, it consistently finds better results across all instances in a shorter time than without them.
Figure 5: The variation of kernel sizes when ablating different reductions. A cell value indicates the difference in kernel size between without using reductions and with using reductions. A higher cell value indicates a larger effect.
Figure 6: The variation of pre-processing time for getting a kernel. A cell value indicates the difference in processing time(s) between without using the corresponding reductions and with using the corresponding reductions. A higher cell value indicates a larger effect.
@@ 6.4 Evaluation of Reduction Rules
Figure 5 shows that the impacts on the kernel size vary among these reduction rules. When comparing individual reductions, $R10$ and $R12$ are more powerful than others in decreasing the kernel size. When comparing a group of reductions, the degree-based rules ($R7$$R11$) are the most powerful in decreasing the kernel size. The neighborhood-based rules ($R3$$R6$) are generally more effective than vertex-pair reduction rules ($R12$$R14$) except for Dblp and WikiTalk. Additionally, although $R1$ and $R2$ do not involve vertex removals, they are also powerful in decreasing the kernel size. In particular, incorporating $R2$ into RLS can further reduce more than 3000 vertices for the BerkStan instance.
It is intuitive that ablating a reduction rule can decrease the overall pre-processing time, but may at the cost of resulting in a larger kernel. However, Figure 6 shows that ablating reduction rules $R3$$R6$ or $R7$$R11$ can even increase the pre-processing time. This is due to the fact that other reductions cost more time when these reduction rules are absent. We also observe that for the larger four instances As-Skitter, Berkstan, WikiTalk and cnr-2000, the inclusion of reductions shows a significant improvement over the pre-processing time.
@@ 7 Conclusion and Future Works
In this paper, we proposed a novel approach to the Generalized Independent Set (GIS) problem using data reduction techniques to tackle the complexity of large graphs. We developed 14 reduction rules, categorized into edge transformation, neighborhood-based, low-degree, and vertex-pair reductions, and designed a reduction-driven local search (RLS) algorithm based on them. Experiments on benchmark instances showed that the RLS algorithm outperformed state-of-the-art methods, especially on large graphs with tens of thousands of vertices. Future work includes exploring advanced reduction techniques for improved efficiency and adapting our algorithms to more real-world scenarios like dynamic graphs.
@@ A Technical Proofs
Consider whether $u$ and $v$ are included in a generalized independent set $S$. There are two cases: either $u\notin S$ or $v\notin S$, and $\{u,v\}\subseteq S$. When either $u\notin S$ or $v\notin S$, the edge $(u,v)$ does not appear in $S$, and thus removing $(u,v)$ has no effect on the net benefit $nb(S)$. If both $u\in S$ and $v\in S$, the edge $(u,v)$ has a zero penalty, which also has no effect on the net benefit $nb(S)$. ∎
@@
Assume, for the sake of contradiction, that there exists $I\in MGIS(G)$ that $u\in I$ and $v\in I$. We have $nb(I)-nb(I\setminus\{u\})=w(u)-\sum_{x\in N_{r}(u)\cap S}p(u,x)\leq w(u)-p(u,v)+\sum_{x\in N_{r}(u)}max(0,-p(u,x))=\tilde{w}(u)-p(u,v)<0$. In conclusion, $I\setminus\{u\}$ has a larger net benefit, which contradicts the optimality of $I$. Therefore, $u$ and $v$ cannot be simultaneously included in any $S\in MGIS(G)$, which is equivalent to the case where $p(u,v)=M$. ∎
The proof of R3 directly follows from the Reduction 2 of Lamm et al. (2019).
@@
We will prove that $u$ is in some set in $MGIS(G)$. Assume, for the sake of contradiction, that $u$ is not included in any set in $MGIS(G)$. Let $I$ be an arbitrary set in $MGIS(G)$. Let $S:=N_{p}(u)\cap I$ and $S^{\prime}:=N_{r}(u)\cap I$. Now, we remove $S$ from $I$ and add $u$ to $I$ to obtain $I^{\prime}$. Consider the difference between two net benefits $nb(I^{\prime})-nb(I)\geq w(u)-\sum_{v\in S}\tilde{w}(v)-\sum_{v\in S^{\prime}}p(u,v)\geq 0$. Therefore, $I^{\prime}\in MGIS(G)$, which contradicts the assumption. So $u$ is in at least one set in $MGIS(G)$. And it holds that $\alpha(G)=\alpha(G^{\prime})+w(u)$. ∎
@@
We prove that $u$ is not in any set in $MGIS(G)$. Given an arbitrary $I\in MGIS(G)$, there are two cases. When none of the neighbors of $u$ is included in $I$, $I$ does not include $u$ because $w(u)<0$. When any subset of neighbors of $u$ is included in $I$, $I$ also does not include $u$ because $w(u)-\sum_{v\in S}p(u,v)<0$ for any subset $S\subseteq N(u)$. Since $I$ is an arbitrary set in $MGIS(G)$, $u$ will not appear in any $I\in MGIS(G)$ and we can directly remove $u$ from $G$. ∎
@@
Observe that there exists a vertex set $I\in MGIS(G)$ that $I\cap N_{p}[u]\neq\varnothing$, because if $I\cap N_{p}[u]=\varnothing$ then $nb(I\cup\{u\})=nb(I)+w(u)\geq nb(I)$ implies $I\cup\{u\}\in MGIS(G)$.
Whenever a vertex $x$ in $N_{p}(u)$ is included in $I$, other vertices in the clique $N_{p}[u]\setminus\{x\}$ must not be included in $I$. Then $I\cup\{u\}\setminus\{x\}\in MGIS(G)$ because $nb(I\cup\{u\}\setminus\{x\})\geq nb(I)+w(u)-w(x)+\sum_{v\in N_{r}(x)}p(v,x)-\sum_{v\in N_{r}(u)}max(0,p(u,v))\geq nb(I)+w(u)-\sum_{v\in N_{r}(u)}max(0,p(u,v))-\max(0,\tilde{w}(x))\geq nb(I)$. In conclusion, $u$ is in some set in $MGIS(G)$. ∎

### v2

@@ 6.4 Evaluation of Reduction Rules
Instance $|V|$ $|V_{ker}|$ CPLEX LA-B&B RLS
without reduction with reduction without reduction with reduction without reduction with reduction
BKV htime(s) BKV htime(s) BKV htime(s) BKV htime(s) BKV htime(s) BKV htime(s)
bio-yeast 1,458 0 68,574 <<1 68,574 <<1 66,545 2 68,574 <<1 68,574 1 68,574 <<1
soc-wiki-Vote 889 0 37,358 <1<1 37,358 <<1 35,973 <<1 37,358 <<1 37,358 1 37,358 <<1
vc-exact_131 2,980 0 109,641 <1<1 109,641 <<1 107,812 14 109,641 <<1 109,567 9 109,641 <<1
web-edu 3,031 98 114,140 <1<1 114,140 <1<1 109,856 10 114,140 21 114,080 12 114,140 <<1
MANN_a81 3,321 0 117,327 <1<1 117,327 <<1 114,616 16 117,327 <<1 117,275 14 117,327 <<1
tech-routers-rf 2,113 0 97,799 <1<1 97,799 <<1 93,690 54 97,799 <<1 97,799 1 97,799 <<1
vc-exact_039 6,795 0 259,542 <1<1 259,542 <<1 254,941 253 259,542 <<1 259,475 13 259,542 <<1
ca-GrQc 4,158 282 174,782 9 174,782 6 167,846 45 174,525 2,998 174,768 7 174,782 11
vc-exact_038 786 697 12,015 592 12,015 585 10,406 1,994 10,782 28 12,015 13 12,015 4
vc-exact_078 11,349 0 438,671 <1<1 438,671 1 431,371 1,074 438,671 1 438,306 16 438,671 <<1
vc-exact_087 13,590 0 518,500 2 518,500 <<1 509,734 2,319 518,500 <<1 518,069 15 518,500 <<1
vc-exact_107 6,402 0 407,878 <1<1 407,878 <<1 396,456 10,646 407,878 <<1 406,379 14 407,878 <<1
vc-exact_151 15,783 0 530,138 2 530,138 <<1 516,445 2,402 530,138 <<1 529,003 18 530,138 <<1
vc-exact_167 15,783 0 465,178 1 465,178 <<1 449,579 10,469 465,178 <<1 463,446 12 465,178 <<1
bio-dmela 7,393 44 302,992 <1<1 302,992 <1<1 280,897 170 302,992 <1<1 302,793 18 302,992 <<1
vc-exact_011 9,877 172 303,793 1 303,793 <1<1 286,137 4,818 303,589 1271 303,352 9 303,793 <<1
web-spam 4,767 487 190,507 5 190,507 <1<1 180,446 53 189,897 <1<1 190,460 18 190,460 13
vc-exact_026 6,140 300 201,509 1 201,509 <1<1 177,902 668 200,868 <1<1 201,458 13 201,503 <<1
vc-exact_001 6,160 134 215,941 2 215,941 <1<1 194,984 9,575 215,941 2 215,934 13 215,941 <<1
vc-exact_024 7,620 1,255 236,109 10,800 236,109 10,800 226,540 244 234,933 11 235,911 14 236,018 12
C1000.9 1,000 1,000 5,263 10,800 5,263 10,800 6,104 430 6,104 430 7,235 16 7,237 10
tech-WHOIS 7,476 338 325,584 5 325,584 <1<1 309,903 218 325,119 255 325,480 16 325,581 13
p-hat700-3 700 700 4,025 10,800 4,025 10,800 3,030 7,436 3,030 7,436 4,473 1 4,473 <<1
vc-exact_008 7,537 574 247,091 3 247,091 <1<1 210,126 9,093 245,289 6,627 247,036 16 247,091 <<1
keller5 776 776 2,373 10,800 2,373 10,800 2,860 6,327 2,860 6,327 3,877 2 3,877 2
vc-exact_194 1,150 1,150 3,734 10,800 3,734 10,800 4,386 1,324 4,386 1,324 5,419 2 5,419 <<1
hamming10-4 1,024 1,024 3,256 10,800 3,256 10,800 3,073 2,069 3,073 2,069 3,660 1 3,660 <<1
brock800_2 800 800 1,177 10,800 1,177 10,800 1,996 10 1,996 10 2,473 1 2,473 <<1
brock800_4 800 800 1,583 10,800 1,583 10,800 1,979 7,215 1,979 7,215 2,501 1 2,501 <<1
ca-HepPh 11,204 2,165 372,096 10,800 372,229 10,800 348,738 422 368,637 2 372,079 14 372,079 17
p-hat700-2 700 700 2,870 10,800 2,870 10,800 2,345 2,078 2,345 2,078 2,971 1 2,971 <<1
vc-exact_196 1,534 1,534 3,916 10,800 3,916 10,800 4,273 66 4,273 66 5,321 2 5,321 <<1
p-hat700-1 700 700 730 10,800 730 10,800 844 5 844 5 844 1 844 <<1
ca-AstroPh 17,903 4,985 604,288 10,800 604,342 10,800 558,736 2,544 593,899 37 603,732 15 603,945 14
C2000.9 2,000 2,000 3,958 10,800 3,958 10,800 5,477 11 5,477 11 6,396 3 6,396 3
DSJC1000.5 1,000 1,000 925 10,800 925 10,800 1,515 694 1,515 694 1,861 1 1,861 <<1
socfb-MIT 6,402 5,494 7,098 10,800 118,114 10,800 110,677 20 113,623 10 129,831 16 129,872 16
p-hat1500-3 1,500 1,500 4,097 10,800 4,097 10,800 3,877 704 3,877 704 7,500 15 7,501 16
socfb-UCSB37 14,917 13,655 16,708 10,800 57,084 10,800 219,965 256 223,780 150 255,887 21 255,898 22
socfb-Duke14 9,885 9,003 11,565 10,800 127,145 10,800 125,940 39 127,269 2,231 151,321 14 151,326 18
p-hat1500-2 1,500 1,500 3,542 10,800 3,542 10,800 2,611 2,150 2,611 2,150 4,253 1 4,253 <<1
socfb-Stanford3 11,586 9,380 7,250 10,800 127,708 10,800 215,381 265 223,398 7,756 251,788 19 251,833 14
socfb-UConn 17,206 16,025 16,996 10,800 51,328 10,800 231,718 322 235,104 240 275,006 16 275,034 20
p-hat1500-1 1,500 1,500 0 10,800 0 10,800 1,184 5,499 1,184 5,499 1,287 2 1,287 1
C2000.5 2,000 2,000 0 10,800 0 10,800 1,581 865 1,581 865 1,849 5 1,849 6
keller6 3,361 3,361 0 10,800 0 10,800 3,545 1,561 3,545 1,561 5,229 19 5,236 16
C4000.5 4,000 4,000 0 10,800 0 10,800 1,611 1,659 1,611 1,659 1,951 15 1,951 22
foodweb-wet 128 125 2,464 <1<1 2,464 <1<1 2,464 8 2,464 6 2,464 <<1 2,464 <<1
foodweb-dry 128 125 2,226 <1<1 2,226 <1<1 2,226 8 2,226 8 2,226 <<1 2,226 <<1
USAir97 332 0 10,308 <1<1 10,308 <1<1 9,507 407 10,308 <1<1 10,308 <<1 10,308 <<1
powergrid 4,941 0 193,339 <1<1 193,339 <1<1 190,537 108 193,339 <1<1 193,194 52 193,339 <<1
CondMat 23,133 1,648 711,517 12 711,517 <1<1 685,442 5,749 708,907 84 709,927 54 711,334 45
Email 265,009 0 13,054,427 118 13,054,427 1 OOM 13,054,427 1 13,054,390 69 13,054,427 1
Epinion 75,879 253 3,052,059 60 3,052,059 1 OOM 3,051,256 196 3,050,415 67 3,052,059 1
Dblp 317,080 11,851 9,737,911 790 9,737,911 39 OOM 9,715,860 307 9,577,609 46 9,734,653 64
cnr-2000 325,557 59,052 13,367,460 10,800 13,641,278 10,800 OOM 11,347,241 10,796 13,361,167 84 13,368,670 482
WikiTalk 2,394,385 8 118,708,326 3,566 118,708,326 333 OOM 118,708,326 333 118,672,055 56 118,708,326 333
BerkStan 685,230 114,191 26,036,210 10,800 26,055,767 10,800 OOM 22,591,394 1,031 25,813,846 200 26,031,704 1,112
As-Skitter 1,696,415 31,173 66,603,360 10,800 66,830,857 10,800 OOM 65,907,243 10,512 59,886,304 200 66,582,254 723
soc-pokec 1,632,803 881,319 4,146,172 10,800 27,037,851 10,800 OOM OOM 36,540,838 350 43,999,402 934
LiveJ 4,846,609 401,307 0 10,800 163,061,840 10,800 OOM OOM 148,758,359 318 168,431,056 904
uk2002 18,483,186 4,806,660 0 100,000 567,666,923 100,000 OOM OOM 633,365,020 1,806 687,089,901 7,678
# of best / # of instances 31/62 36/62 3/62 21/62 22/62 51/62
average running time(s) 7,096.5 7,028.8 2,007.5 1,441.1 61.3 202.6
Comparative results with vs. without reduction. This experiment evaluates algorithms with and without data reduction. For LA-B&B and CPLEX with the reductions, we used kernel graphs as input. For the RLS without reduction, all reductions were skipped, and random initialization was applied. ALS cannot handle kernel graphs due to unsupported negative vertex profits and edge penalties.
The comparative results are shown in Table 5. In general, the reduction rules can help all algorithms to find better solutions. Specifically, with the reduction techniques, LA-B&B is able to solve some instances which it cannot solve without reduction, e.g., WikiTalk (1.6 million vertices, 11.1 million edges). For some large instances where original LA-B&B run out-of-memory, the new LA-B&B with the reductions can report a feasible solution or solve them to optimality within the cutoff time. With the reduction techniques, CPLEX can solve some instances with less time. For two large instances LiveJ and uk2002 where CPLEX fails, CPLEX with the reductions is able to report feasible solutions. As for the RLS, with these reduction rules, it consistently finds better results across all instances in a shorter time than without them.
Figure 5: Variation of kernel sizes when ablating different reductions. Each cell shows the difference in kernel size between not using and using the corresponding reductions; higher values indicate larger effects.
Figure 6: Variation of pre-processing time for getting a kernel. Each cell shows the difference in time (seconds) between not using and using the corresponding reductions; higher values indicate larger effects.
@@
Figure 5 shows that the impacts on the kernel size vary among these reduction rules. When comparing individual reductions, $R10$ and $R12$ are more powerful than the other reduction rules in decreasing the kernel size. When comparing a group of reductions, the degree-based rules (R7R7R11R11) are the most powerful in decreasing the kernel size. The neighborhood-based rules (R3R3R6R6) are generally more effective than the vertex-pair reduction rules (R12R12R14R14) except for Dblp and WikiTalk. Additionally, although $R1$ and $R2$ do not involve vertex removals, they are also powerful in decreasing the kernel size. In particular, incorporating $R2$ into the RLS can further reduce more than 3,000 vertices for the BerkStan instance.
Intuitively, ablating a reduction rule can decrease the overall pre-processing time, but may at the cost of resulting in a larger kernel. However, Figure 6 shows that ablating reduction rules R3R3R6R6 or R7R7R11R11 can even increase the pre-processing time. This occurs because the remaining reductions become more time-consuming in the absence of these rules. We also observe that for the four larger instances As-Skitter, Berkstan, WikiTalk and cnr-2000, the inclusion of reductions shows a significant improvement over the pre-processing time.
@@ 7 Conclusion and Future Works
In this paper, we proposed a novel approach to the Generalized Independent Set (GIS) problem using data reduction techniques to tackle the complexity of large graphs. We developed 14 reduction rules, categorized into edge transformation, neighborhood-based, low-degree, and vertex-pair reductions, and designed a reduction-driven local search (RLS) algorithm based on them. Experiments on benchmark instances showed that the RLS algorithm outperformed state-of-the-art methods, especially on large graphs with tens of thousands of vertices. Future work will focus on developing more advanced reduction techniques to further improve efficiency and extending our algorithms to real-world scenarios such as dynamic graphs with changing weights.
@@ A Technical Proofs
Consider whether $u$ and $v$ are included in a generalized independent set $S$. There are four cases: $u\in S$ and $v\in S$, $u\in S$ and $v\notin S$, $u\notin S$ and $v\in S$, and $u\notin S$ and $v\notin S$. When either $u\notin S$ or $v\notin S$, the edge $e=\{u,v\}$ does not exist in $G[S]$, and thus removing $e$ has no effect on the net benefit $nb(S)$. If both $u\in S$ and $v\in S$, the edge $\{u,v\}$ has a zero penalty, which has no effect on the net benefit $nb(S)$ either. ∎
@@
Assume, for the sake of contradiction, that there exists $I\in MGIS(G)$ such that $u\in I$ and $v\in I$. Assume, without loss of generality, that $\tilde{w}(u)\leq\tilde{w}(v)$. We have $nb(I)-nb(I\setminus\{u\})=w(u)-\sum_{x\in N_{r}(u)\cap{I}}p(\{u,x\})\leq w(u)-p(\{u,v\})+\sum_{x\in N_{r}(u)}max(0,-p(\{u,x\}))=\tilde{w}(u)-p(\{u,v\})<0$. In conclusion, $I\setminus\{u\}$ has a larger net benefit without containing permanent edges, which contradicts the optimality of $I$. Therefore, $u$ and $v$ cannot be simultaneously included in any $S\in MGIS(G)$, which is equivalent to the case where $\{u,v\}\in E_{p}$. ∎
Proof of R3.
We prove that $u$ is in at least one set in $MGIS(G)$. Assume, for the sake of contradiction, that $u$ is not included in any set in $MGIS(G)$. Let $I$ be an arbitrary set in $MGIS(G)$. It is clear that at least one vertex in $N(u)$ must be included in $I$. Otherwise, $I\cup\{u\}$ implies with a larger net benefit without including any permanent edge, a contradiction, since $w(u)\geq w^{+}(N(u))$, $nb(I\cup\{u\}\setminus N(u))\geq nb(I)$. So $u$ is in at least one set in $MGIS(G)$. It holds that $\alpha(G)=\alpha(G^{\prime})+w(u)$. ∎
@@
We prove that $u$ is in at least one set in $MGIS(G)$. Assume, for the sake of contradiction, that $u$ is not included in any set in $MGIS(G)$. Let $I$ be an arbitrary set in $MGIS(G)$. Let $S:=N_{p}(u)\cap I$ and $S^{\prime}:=N_{r}(u)\cap I$. Now, we remove $S$ from $I$ and add $u$ to $I$ to obtain $I^{\prime}$. Consider the difference between two net benefits $nb(I^{\prime})-nb(I)\geq w(u)-\sum_{v\in S}\tilde{w}(v)-\sum_{v\in S^{\prime}}p(\{u,v\})\geq 0$. Moreover, $G[I^{\prime}]$ does not contain permanent edges. Therefore, $I^{\prime}\in MGIS(G)$, which contradicts the assumption. So $u$ is in at least one set in $MGIS(G)$. It holds that $\alpha(G)=\alpha(G^{\prime})+w(u)$. ∎
@@
We prove that $u$ is not in any set in $MGIS(G)$. Given an arbitrary $I\in MGIS(G)$, there are two cases. When none of the neighbors of $u$ is included in $I$, II does not include $u$ because $w(u)<0$. When any subset of neighbors of $u$ is included in $I$, II does not include $u$ either because $w(u)-\sum_{v\in S}p(\{u,v\})<0$ for any subset $S\subseteq N(u)$. Since $I$ is an arbitrary set in $MGIS(G)$, $u$ will not appear in any $I\in MGIS(G)$ and we can directly remove $u$ from $G$. ∎
@@
Observe that there exists a vertex set $I\in MGIS(G)$ such that $I\cap N_{p}[u]\neq\varnothing$, because if $I\cap N_{p}[u]=\varnothing$ then $nb(I\cup\{u\})=nb(I)+w(u)\geq nb(I)$ implies $I\cup\{u\}\in MGIS(G)$.
Whenever a vertex $x$ in $N_{p}(u)$ is included in $I$, other vertices in the clique $N_{p}[u]\setminus\{x\}$ must not be included in $I$. Then $I\cup\{u\}\setminus\{x\}\in MGIS(G)$ because $G[I\cup\{u\}\setminus\{x\}]$ does not contain permanent edges, and $nb(I\cup\{u\}\setminus\{x\})\geq nb(I)+w(u)-w(x)+\sum_{v\in N_{r}(x)}p(\{v,x\})-\sum_{v\in N_{r}(u)}max(0,p(\{u,v\}))\geq nb(I)+w(u)-\sum_{v\in N_{r}(u)}max(0,p(\{u,v\}))-\max(0,\tilde{w}(x))\geq nb(I)$. In conclusion, $u$ is in at least one set in $MGIS(G)$. ∎